OpenStudy (anonymous):

if there are 60 students, how many 6th graders own cats

OpenStudy (anonymous):

i will give u the picture

OpenStudy (anonymous):

35% of 60

OpenStudy (anonymous):

21 6th graders will have cats.

OpenStudy (anonymous):

with an example and expination

OpenStudy (anonymous):

\[.35\times 60\]

OpenStudy (anonymous):

\[60*0.30= answer.\]

OpenStudy (anonymous):

AND I WILL GIVE A MEADEL

OpenStudy (anonymous):

err 35.

OpenStudy (anonymous):

explain and an example

OpenStudy (cwrw238):

you don't own cats - they own you!

OpenStudy (cwrw238):

0.35 * 60 = 10*0.35 * 6 = 3.5 * 6 = ?

OpenStudy (anonymous):

21

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw